OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2).

OpenStudy (anonymous):

A line perpendicular to this will have a slope -(1/5) y = -x/5 + b it will pass through x = -3, y=2. Find b.